n=50, p = 0.02, x=2 P(2) = _______

A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. 

 n=50, p = 0.02, x=2 


P(2) = _______ 
Answer #1

Solution :
Given that ,
p = 0.02
1 - p = 1 - 0.02 = 0.98
n = 50
Using binomial probability formula ,
P(X = 2) = ((50! / 2! (50 - 2)!) * 0.022 * (0.98)50 - 2
P(2) = ((50! / 2! (48)!) * 0.022 * (0.98)48 = 0.1858
Probability = 0.1858